OpenStudy (anonymous):

Ok people i have the worlds hardest math question literary i dont need help on it but i want 2 see if anyone can get it its just for fun and i have the answer ok so here it is: 3^2 + 4^2 = 5^2 and 5^2 + 12^2 = 13^2

OpenStudy (anonymous):

YA ITS TRUE

OpenStudy (anonymous):

what's the question?

OpenStudy (anonymous):

First one: True Second one: True. They are both true.

OpenStudy (anonymous):

All right... this means...? I mean, you can generate these as: \[ \forall m,n\in\mathbb{Z}\\ \left(m^2-n^2\right)^2+\left(2mn\right)^2=\left(m^2+n^2\right)^2 \]But, your question is not a question, it is a statement.

OpenStudy (anonymous):

Fermat claimed that if n > 2 the equation x^n + y^n = z^n has no solutions with x,y,z positive integers.

OpenStudy (anonymous):

its basicly a true or false question

OpenStudy (anonymous):

Yeah, I agree with @LolWolf . It's asking us if it's true or false..

OpenStudy (anonymous):

3^2 + 4^2 = 5^2 and 5^2 + 12^2 = 13^2. Fermat claimed that if n > 2 the equation x^n + y^n = z^n has no solutions with x,y,z positive integers. This was finally proved in 1994 after over 300 years. But the methods used were totally indirect. Namely by proving something really important which had been shown a few years earlier to have Fermat's last theorem as a corollary. (But the proof of this big theorem was about 200 pages long & had to be checked extremely carefully by a dozen experts. Since then it has been shortened by others using newer ideas, but still can't be comprehended by almost all professional mathematicians, One has to believe the 100 -200 who really do understand it. NOT satisfactory.) It's fair to say that Fermat's last theorem is merely a very hard puzzle, which became a major challenge because lots of very smart people failed to prove it. Gauss, one of the greatest mathematicians of all time, thought FLT was uninteresting. He said one could easily make up similar problems just as hard. There are 7 "Millennium Problems", proposed in 2000, which are extremely hard & important. For what it's worth, there's a million dollar prize for each. One has been solved.

OpenStudy (anonymous):

Oh, hey, by Wiles proof of the Taniyama–Shimura conjecture, such is true. If we use Ken Ribet's proof of the elipson conjecture, Fermat's last theorem is implied. Q.E.D. :) (Of course, each proof requires like 500 books to understand, but they've already been proven).

OpenStudy (anonymous):

Oh, nice: http://answers.yahoo.com/question/index?qid=20110901225732AAD88Uj Is that really what you looked up?

OpenStudy (anonymous):

pfft yall r so borin jus playin round bein bored and ur like -_- the answer is arrgg lol

OpenStudy (anonymous):

We're mathematicians/engineers, and, like everyone knows, we sit at home staring at a page filled with equations for hours on end without eating or drinking anything, hardly moving. We never go outside, and we only talk in numbers and expressions. We have only one friend, who can comprehend us, but he, himself, is a mathematician or physicist, and does the same. We have no social life, parties don't exist, we think 'fun' is overrated, and life is all just one big equation. Of course, we must be boring. There's no other logical conclusion!

OpenStudy (anonymous):

lol okkie thnn

OpenStudy (anonymous):

Please tell me you didn't just believe what I said...

OpenStudy (anonymous):

lmao no

OpenStudy (anonymous):

All right, just making sure, haha.